# A multi-messenger hierarchical triple merger gravitational-wave event pair GW190514-GW190521 inside AGN J124942.3 + 344929

### v2

@@ I Introduction
GW190521 [1] is one of the most massive binary black hole (BBH) merger observed during the first three LIGO-Virgo observing runs, with component masses of approximately $85\,M_{\odot}$ and $66\,M_{\odot}$, producing a remnant black hole (BH) of around $142\,M_{\odot}$, which can be classified as an intermediate mass BH. GW190514 [2] is another BBH merger, with component masses of approximately $39\,M_{\odot}$ and $28\,M_{\odot}$, resulting in remnant BH of around $67\,M_{\odot}$, which overlaps with the second mass range of GW190521. Moreover, there is significant spatial overlap between the sky localizations and distances of GW190514 and GW190521 (see Fig. 1), suggesting a plausible scenario where these two events are part of a hierarchical triple merger. In this scenario, GW190514 would represent the first merger, producing a remnant BH that later paired with another BH to form the hierarchical system responsible for GW190521 [3].
In addition, an optical electromagnetic (EM) counterpart candidate, ZTF19abanrhr, was reported by Graham et al. [4], detected by the Zwicky Transient Facility approximately 34 days after the gravitational wave (GW) event GW190521. Evidence in favor of an association between the GW event and the EM counterpart was also reported in [5, 6] (although [7, 8]) This counterpart, confirmed as a flare within the active galactic nuclei (AGN) AGN J124942.3 + 344929, was at the 78% spatial contour for GW190521’s sky localization. Interestingly, GW190514’s 90% sky localization also coincides with the position of ZTF19abanrhr (see Fig. 1). This raises an intriguing question: Could both GW190514 and GW190521 be a hierarchical triple merger originating from AGN J124942.3 + 344929, with ZTF19abanrhr serving as their EM counterpart?
@@ II Methods
We explore the potential association between GW190514 and GW190521 as hierarchical triple merger in AGN J124942.3 + 344929, associated with ZTF19abanrhr, using data ($d_{1}$ and $d_{2}$ for the gravitational signals [^1] https://zenodo.org/record/5117702, and $s$ for the EM signal [^2] https://lasair-ztf.lsst.ac.uk/objects/ZTF19abanrhr/) from the GWTC-2.1 release [2] and ZTF [20]. This is a hypothesis-testing question. Our analysis employs a Bayesian criterion to assess whether signals observed from these separate data sets share a common origin. The association ($\mathcal{H}_{1}$) at least requires that i) the localizations of GW190514, GW190521, and ZTF19abanrhr must be consistent (see Fig. 1); and ii) the final mass of GW190514 should align with the second mass in GW190521. In contrary, the coincidence ($\mathcal{H}_{0}$) corresponds to the scenario where the three detections are associated by chance. As a result, the odds $\mathcal{O}^{1}_{0}$ between the hypotheses $\mathcal{H}_{1}$ and $\mathcal{H}_{0}$ is given by: $\mathcal{O}^{1}_{0}(d_{1},d_{2},s)=\mathcal{P}^{1}_{0}\,\mathcal{B}^{1}_{0}(d_{1},d_{2},s)$, where $\mathcal{P}^{1}_{0}$ is the prior odds; and $\mathcal{B}^{1}_{0}(d_{1},d_{2},s)=p(d_{1},d_{2},s|\mathcal{H}_{1})/p(d_{1},d_{2},s|\mathcal{H}_{0})$ is the Bayes factor defined by the ratio of the evidences of the hypotheses.
The derivations and computational framework presented as follows build upon the foundational methodology established in Morton et al. [5], who demonstrated the astrophysical association between GW190521 and flare ZTF19abanrhr as its EM counterpart. Here we investigate the association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr.
@@ II.1 Environmental effects
The first effect, related to the gravitational potential of the supermassive BH, leads to gravitational redshift.
@@
where $R_{\rm s}$ represents the Schwarzschild radius of the supermassive BH, and $r$ is the distance between the BBH and the supermassive BH. The gravitational redshift is calculated assuming a non-spinning supermassive BH, although supermassive BHs in AGN disks are typically expected to spin [23]. This assumption holds because the effect of spin becomes negligible at the distances relevant for BBHs with respect to the supermassive BH [24, 25, 26].
@@
In this case, the detected mass of a redshifted source differs from intrinsic mass in the source frame by a factor equal to the combined effects of all relevant redshifts, including gravitational and relativistic redshifts [32],
@@ II.2 Bayesian criterion
In order to assess whether it is more likely that the two GW events, GW190514 and GW190521 (as a hierarchical triple merger), as well as the EM transient ZTF19abanrhr, originate from a common source, or whether the three detections are associated by chance, we employ a Bayesian statistical analysis. Given three detections in the GW190514 data set $d_{1}$, the GW190521 data set $d_{2}$, and the ZTF19abanrhr data set $s$, the odds $\mathcal{O}^{1}_{0}$ between the association model $\mathcal{H}_{1}$ and the coincidence model $\mathcal{H}_{0}$ is given by:
@@
is the Bayes factor and $\mathcal{P}^{1}_{0}=p(\mathcal{H}_{1})/p(\mathcal{H}_{0})$ is the prior odds. The prior odds for the two models can be expressed as the product of two components: (1) the odds that the EM transient is associated with one of the two GW events, estimated to be $\sim 1/13$ [8], based on the fact that approximately 13 transients similar to ZTF19abanrhr were detected in the Zwicky Transient Facility alert stream over the relevant observing period [4]; (2) the odds that both mergers in the hierarchical triple system are found in the observed BBHs, estimated to be $\sim 1/1000$ [34, 35], corresponding to the expected fraction of hierarchical triple mergers among detectable events in optimistic scenarios. The resulting combined astrophysical prior odds is approximately $1/13{,}000$. For robustness, we adopt a slightly more conservative value of $10^{-5}$ in our analysis, noting that the $1/1000$ estimate represents an optimistic upper limit and the typical value is likely smaller. This choice (approximately eight times smaller) ensures that the preference for the proposed association remains significant even under stricter prior assumptions.
The Bayes factor is the ratio of the evidences for the two hypothetical models. The evidence for a model can be computed via the marginalization over all parameters $\vec{\theta}$ required by the model $\mathcal{H}_{i}$,
@@
The first term of equation (9), $p(d_{1},d_{2},s|D_{L},\Omega,m^{z},\mathcal{H}_{i})$ can be written as
@@
The second term in equation (9), $p(D_{L},\Omega,m^{z}|\mathcal{H}_{i})$, is the prior conditioned under the hypothetical model $\mathcal{H}_{i}$. In general, these parameters are a priori independent, we can factorize it as
@@
Substituting equations (10)&(11) into equation (9):
@@ II.3 Association model
In the association model $\mathcal{H}_{1}$, the GW events GW190514 and GW190521, as a hierarchical triple merger, and the flare ZTF19abanrh share a common origin. The localization is fixed to the host AGN J124942.3 + 344929, and as a result, the observed mass and distance of the source are altered due to environmental effects. In this case, the GW posterior probability densities used in our work are the detector-frame effective primary mass $m^{z,{\rm eff}}$ and effective luminosity distance $D_{L}^{\rm eff}$. These two parameters are independently measured in GW detection; however, under the common origin hypothesis, they depend on the additional redshifts introduced by environmental effects.
@@
the redshift $z_{\rm AGN}=0.438$ of AGN J124942.3 + 344929333https://skyserver.sdss.org/dr12/en/tools/explore/Summary.aspx?id=1237665128546631763. The expressions for $z_{\rm grav}(r)$ and $z_{\rm rel}(r,\phi)$ in equation (14) are the ones in equations (1) and (2), respectively. Here we adopt the prior $p(\phi)$ to be uniform in ${\rm cos}(\phi)$. The radial distance prior, $p(r)$, between the BBH and the supermassive BH is predicted based on migration traps in AGN disks [36]. We assume that this radial distance is uniformly distributed between 24.5 and 331 Schwarzschild radii [36]. It is important to note that the precise location of migration traps can vary depending on the disk model, and in some cases, there may not be a migration trap at all [37, 38, 39]. However, our results are robust and are not significantly affected by these uncertainties [5].
In the same way, the prior conditioned under the hypothetical model $\mathcal{H}_{1}$ reads $p(m^{z}|\mathcal{H}_{1})\rightarrow p(m^{z,{\rm eff}}|\mathcal{H}_{1})$:
@@
where the cosmological redshift $z_{c}\equiv z_{\rm AGN}$ and $p(m^{z}|\mathcal{H}_{1})=(1+z_{\rm c})\,p(m|\mathcal{H}_{1})$. The astrophysical prior $p(m|\mathcal{H}_{1})$ represents the mass distribution of BBHs formed in AGN disks, which is presented in the Appendix II.5.
@@ II.4 Coincidence model
In the coincidence model $\mathcal{H}_{0}$, the three transients GW190514, GW190521, and ZTF19abanrhr are unrelated and happen to appear within the same localization by chance. In this scenario, there are no environmental effects from AGNs, and therefore, the only redshift present is the cosmological one. As a result, the prior conditioned under the hypothetical model $\mathcal{H}_{0}$ is given by,
@@ II.6 Models including primary masses
where $m^{\prime,z}$ and $m^{\prime\prime,z}$ are the primary masses of GW190514 and GW190521 in the detector frame, respectively; $p(m^{\prime,z}|\mathcal{H}_{1})$ and $p(m^{\prime\prime,z}|\mathcal{H}_{1})$ are the priors used during the parameter estimation run; and $p^{\rm 1g}_{m_{\rm 1}}(m^{\prime,z}|\mathcal{H}_{1})$ and $p^{\rm hier}_{m_{\rm 1}}(m^{\prime\prime,z}|\mathcal{H}_{1})$ ($p^{\rm 1g}_{m_{\rm 1}}(m^{\prime,z}|\mathcal{H}_{0})$ and $p^{\rm hier}_{m_{\rm 1}}(m^{\prime\prime,z}|\mathcal{H}_{0})$) are the priors conditioned of primary detector-frame masses of first generation and hierarchical mergers in the AGN formation channel (in formation channels other than AGN disks), respectively.
@@ III Results
Furthermore, we find that the primary masses of these two GW events are also crucial for enhancing the probability of the association compared to the coincidence [5]. This is because BBH mergers in AGN disks are generally expected to have more massive components compared to other channels, such as isolated binary evolution or star clusters, due to accretion and frequent hierarchical mergers [9, 10, 11]. Therefore, we also consider the primary masses ($m_{1}$) of the GW events to assess the multi-messenger coincidence significance for the three transients (see Section II.6). Figure 3 presents the posterior probability densities of the primary masses of the two GW events, comparing the distributions from the GWTC-2.1 [2] with those from the association model (labeled as ‘Association with $m_{1}$’, green). As shown in this figure, the association model with $m_{1}$ provides better constraints, slightly favoring a more massive BH for GW190521 and a less massive one for GW190514, compared to the GWTC-2.1 distributions [2]. Notably, GW190521’s primary BH mass falls within the mass gap with nearly 100% probability, strongly suggesting that it is a hierarchical merger [48, 49, 50, 51, 52, 53, 54, 55], potentially a second (or higher) generation merger resulting from GW190514 in an AGN environment. Incorporating the primary mass information significantly strengthens the association hypothesis, increasing the Bayes factor by approximately two orders of magnitude, corresponding to a log Bayes factor of ${\rm log}\mathcal{B}=21.5$ and an odds ratio of $\mathcal{O}\sim$$12\times 10^{4}:$ 1. This result aligns with the expectation that more massive BBH mergers are more likely to occur in the AGN formation channel [9, 10, 11]. Additionally, we find that the inclusion of primary mass data allows slightly for better constraints on the associated mass between GW190514 and GW190521 (labeled as ‘Association with $m_{1}$’, green, in Fig. 2).
@@ IV Conclusions and discussion
In this work, we investigate the potential association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr, taking into account of sky position, distance, and mass of the sources using a Bayesian criterion. Although the previous study has shown that GW190514 and GW190521 are a plausible hierarchical triple merger, their analysis is not significant with $p$-value of $\sim$0.14 [3] (generally, $p$-value of $\leq$0.05 is considered ‘significant’). Here, we introduce a new approach and propose using multi-messenger [4] signals to analyze the probability of the GW190514-GW190521 hierarchical triple merger. Our analysis (considering the primary masses) results greatly increase the probability, finding that a log Bayes factor of ${\rm log}\mathcal{B}=16.8~{}(21.5)$, corresponding to an odds ratio of $\mathcal{O}\sim$$1199~{}(2\times 10^{4}):$ 1, assuming an astrophysical prior odds of $\mathcal{P}\sim 10^{-5}$. This suggest strong evidence for the first hierarchical triple merger associated with an EM counterpart in the AGN formation channel. The finding of the multi-messenger hierarchical triple merger in the AGN is crucial for understanding the formation mechanisms of massive BHs, the role of AGNs in hierarchical mergers, and the implications of multi-messenger astronomy

### v3

# Multimessenger hierarchical triple merger gravitational-wave event pair GW190514-GW190521 inside AGN J124942.3 + 344929

@@ I Introduction
GW190521 [1] is one of the most massive binary black hole (BBH) mergers observed during the first three LIGO-Virgo observing runs, with component masses of approximately $85\,M_{\odot}$ and $66\,M_{\odot}$, producing a remnant black hole (BH) of around $142\,M_{\odot}$, which can be classified as an intermediate mass BH. GW190514 [2] is another BBH merger, with component masses of approximately $39\,M_{\odot}$ and $28\,M_{\odot}$, resulting in a remnant BH of around $67\,M_{\odot}$, which overlaps with the second mass range of GW190521. Moreover, there is significant spatial overlap between the sky localizations and distances of GW190514 and GW190521 (see Fig. 1), suggesting a plausible scenario where these two events are part of a hierarchical triple merger. In this scenario, GW190514 would represent the first merger, producing a remnant BH that later paired with another BH to form the hierarchical system responsible for GW190521 [3].
In addition, an optical electromagnetic (EM) counterpart candidate, ZTF19abanrhr, was reported by Graham et al. [4], detected by the Zwicky Transient Facility approximately 34 days after the gravitational-wave (GW) event GW190521. Evidence in favor of an association between the GW event and the EM counterpart was also reported in [5, 6] (although see [7, 8]). This counterpart, confirmed as a flare within the active galactic nuclei (AGN) AGN J124942.3 + 344929, was at the 78% spatial contour for GW190521’s sky localization. Interestingly, GW190514’s 90% sky localization also coincides with the position of ZTF19abanrhr (see Fig. 1). This raises an intriguing question: Could both GW190514 and GW190521 be a hierarchical triple merger originating from AGN J124942.3 + 344929, with ZTF19abanrhr serving as their EM counterpart?
@@ II Methods
We explore the potential association between GW190514 and GW190521 as a hierarchical triple merger in AGN J124942.3 + 344929, associated with ZTF19abanrhr, using data ($d_{1}$ and $d_{2}$ for the gravitational signals [^1] https://zenodo.org/record/5117702, and $s$ for the EM signal [^2] https://lasair-ztf.lsst.ac.uk/objects/ZTF19abanrhr/) from the GWTC-2.1 release [2] and ZTF [20]. This is a hypothesis-testing question. Our analysis employs a Bayesian criterion to assess whether signals observed from these separate datasets share a common origin. The association ($\mathcal{H}_{1}$) at least requires that i) the localizations of GW190514, GW190521, and ZTF19abanrhr must be consistent (see Fig. 1); and ii) the final mass of GW190514 should align with the second mass in GW190521. In contrary, the coincidence ($\mathcal{H}_{0}$) corresponds to the scenario where the three detections are associated by chance. As a result, the odds $\mathcal{O}^{1}_{0}$ between the hypotheses $\mathcal{H}_{1}$ and $\mathcal{H}_{0}$ is given by: $\mathcal{O}^{1}_{0}(d_{1},d_{2},s)=\mathcal{P}^{1}_{0}\,\mathcal{B}^{1}_{0}(d_{1},d_{2},s)$, where $\mathcal{P}^{1}_{0}$ is the prior odds; and $\mathcal{B}^{1}_{0}(d_{1},d_{2},s)=p(d_{1},d_{2},s|\mathcal{H}_{1})/p(d_{1},d_{2},s|\mathcal{H}_{0})$ is the Bayes factor defined by the ratio of the evidences of the hypotheses.
The derivations and computational framework presented as follows build upon the foundational methodology established in Morton et al. [5], who demonstrated the astrophysical association between GW190521 and flare ZTF19abanrhr as its EM counterpart. Here, we investigate the association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr.
@@ II.1 Environmental effects
The first effect, related to the gravitational potential of the supermassive BH, leads to gravitational redshift:
@@
where $R_{\rm s}$ represents the Schwarzschild radius of the supermassive BH, and $r$ is the distance between the BBH and the supermassive BH. The gravitational redshift is calculated assuming a non-spinning supermassive BH, although supermassive BHs in AGN disks are typically expected to spin [23]. This assumption holds, because the effect of spin becomes negligible at the distances relevant for BBHs with respect to the supermassive BH [24, 25, 26].
@@
In this case, the detected mass of a redshifted source differs from intrinsic mass in the source frame by a factor equal to the combined effects of all relevant redshifts, including gravitational and relativistic redshifts [32]:
@@ II.2 Bayesian criterion
In order to assess whether it is more likely that the two GW events, GW190514 and GW190521 (as a hierarchical triple merger), as well as the EM transient ZTF19abanrhr, originate from a common source, or whether the three detections are associated by chance, we employ a Bayesian statistical analysis. Given three detections in the GW190514 dataset $d_{1}$, the GW190521 dataset $d_{2}$, and the ZTF19abanrhr dataset $s$, the odds $\mathcal{O}^{1}_{0}$ between the association model $\mathcal{H}_{1}$ and the coincidence model $\mathcal{H}_{0}$ is given by
@@
is the Bayes factor and $\mathcal{P}^{1}_{0}=p(\mathcal{H}_{1})/p(\mathcal{H}_{0})$ is the prior odds. The prior odds for the two models can be expressed as the product of two components: (i) the odds that the EM transient is associated with one of the two GW events, estimated to be $\sim 1/13$ [8], based on the fact that approximately 13 transients similar to ZTF19abanrhr were detected in the Zwicky Transient Facility alert stream over the relevant observing period [4]; (ii) the odds that both mergers in the hierarchical triple system are found in the observed BBHs, estimated to be $\sim 1/1000$ [34, 35], corresponding to the expected fraction of hierarchical triple mergers among detectable events in optimistic scenarios. The resulting combined astrophysical prior odds is approximately $1/13{,}000$. For robustness, we adopt a slightly more conservative value of $10^{-5}$ in our analysis, noting that the $1/1000$ estimate represents an optimistic upper limit and the typical value is likely smaller. This choice (approximately eight times smaller) ensures that the preference for the proposed association remains significant even under stricter prior assumptions.
The Bayes factor is the ratio of the evidences for the two hypothetical models. The evidence for a model can be computed via the marginalization over all parameters $\vec{\theta}$ required by the model $\mathcal{H}_{i}$:
@@
The first term in Equation (9), $p(d_{1},d_{2},s|D_{L},\Omega,m^{z},\mathcal{H}_{i})$ can be written as
@@
The second term in Equation (9), $p(D_{L},\Omega,m^{z}|\mathcal{H}_{i})$, is the prior conditioned under the hypothetical model $\mathcal{H}_{i}$. In general, these parameters are a priori independent; we can factorize it as
@@
Substituting Equations (10)&(11) into Equation (9)
@@ II.3 Association model
In the association model $\mathcal{H}_{1}$, the GW events GW190514 and GW190521, as a hierarchical triple merger, and the flare ZTF19abanrhr share a common origin. The localization is fixed to the host AGN J124942.3 + 344929, and as a result, the observed mass and distance of the source are altered due to environmental effects. In this case, the GW posterior probability densities used in our work are the detector-frame effective primary mass $m^{z,{\rm eff}}$ and effective luminosity distance $D_{L}^{\rm eff}$. These two parameters are independently measured in GW detection; however, under the common origin hypothesis, they depend on the additional redshifts introduced by environmental effects.
@@
the redshift $z_{\rm AGN}=0.438$ of AGN J124942.3 + 344929333https://skyserver.sdss.org/dr12/en/tools/explore/Summary.aspx?id=1237665128546631763. The expressions for $z_{\rm grav}(r)$ and $z_{\rm rel}(r,\phi)$ in Equation (14) are the ones in Equations (1) and (2), respectively. Here, we adopt the prior $p(\phi)$ to be uniform in ${\rm cos}(\phi)$. The radial distance prior, $p(r)$, between the BBH and the supermassive BH is predicted based on migration traps in AGN disks [36]. We assume that this radial distance is uniformly distributed between 24.5 and 331 Schwarzschild radii [36]. It is important to note that the precise location of migration traps can vary depending on the disk model, and in some cases, there may not be a migration trap at all [37, 38, 39]. However, our results are robust and are not significantly affected by these uncertainties [5].
In the same way, the prior conditioned under the hypothetical model $\mathcal{H}_{1}$ reads $p(m^{z}|\mathcal{H}_{1})\rightarrow p(m^{z,{\rm eff}}|\mathcal{H}_{1})$
@@
where the cosmological redshift $z_{c}\equiv z_{\rm AGN}$ and $p(m^{z}|\mathcal{H}_{1})=(1+z_{\rm c})\,p(m|\mathcal{H}_{1})$. The astrophysical prior $p(m|\mathcal{H}_{1})$ represents the mass distribution of BBHs formed in AGN disks, which is presented in Section II.5.
@@ II.4 Coincidence model
In the coincidence model $\mathcal{H}_{0}$, the three transients GW190514, GW190521, and ZTF19abanrhr are unrelated and happen to appear within the same localization by chance. In this scenario, there are no environmental effects from AGNs, and, therefore, the only redshift present is the cosmological one. As a result, the prior conditioned under the hypothetical model $\mathcal{H}_{0}$ is given by
@@ II.6 Models including primary masses
respectively, where $m^{\prime,z}$ and $m^{\prime\prime,z}$ are the primary masses of GW190514 and GW190521 in the detector frame, respectively; $p(m^{\prime,z}|\mathcal{H}_{1})$ and $p(m^{\prime\prime,z}|\mathcal{H}_{1})$ are the priors used during the parameter estimation run; and $p^{\rm 1g}_{m_{\rm 1}}(m^{\prime,z}|\mathcal{H}_{1})$ and $p^{\rm hier}_{m_{\rm 1}}(m^{\prime\prime,z}|\mathcal{H}_{1})$ ($p^{\rm 1g}_{m_{\rm 1}}(m^{\prime,z}|\mathcal{H}_{0})$ and $p^{\rm hier}_{m_{\rm 1}}(m^{\prime\prime,z}|\mathcal{H}_{0})$) are the priors conditioned of primary detector-frame masses of first generation and hierarchical mergers in the AGN formation channel (in formation channels other than AGN disks), respectively.
@@ III Results
Furthermore, we find that the primary masses of these two GW events are also crucial for enhancing the probability of the association compared to the coincidence [5]. This is because BBH mergers in AGN disks are generally expected to have more massive components compared to other channels, such as isolated binary evolution or star clusters, due to accretion and frequent hierarchical mergers [9, 10, 11]. Therefore, we also consider the primary masses ($m_{1}$) of the GW events to assess the multimessenger coincidence significance for the three transients (see Section II.6). Figure 3 presents the posterior probability densities of the primary masses of the two GW events, comparing the distributions from the GWTC-2.1 [2] with those from the association model (labeled as ‘Association with $m_{1}$’, green). As shown in this figure, the association model with $m_{1}$ provides better constraints, slightly favoring a more massive BH for GW190521 and a less massive one for GW190514, compared to the GWTC-2.1 distributions [2]. Notably, GW190521’s primary BH mass falls within the mass gap with nearly 100% probability, strongly suggesting that it is a hierarchical merger [48, 49, 50, 51, 52, 53, 54, 55], potentially a second (or higher) -generation merger resulting from GW190514 in an AGN environment. Incorporating the primary mass information significantly strengthens the association hypothesis, increasing the Bayes factor by approximately 2 orders of magnitude, corresponding to a log Bayes factor of ${\rm log}\mathcal{B}=21.5$ and an odds ratio of $\mathcal{O}\sim$$12\times 10^{4}:$ 1. This result aligns with the expectation that more massive BBH mergers are more likely to occur in the AGN formation channel [9, 10, 11]. Additionally, we find that the inclusion of primary mass data allows slightly for better constraints on the associated mass between GW190514 and GW190521 (labeled as ‘Association with $m_{1}$’, green, in Fig. 2).
@@ IV Conclusions and discussion
In this work, we investigate the potential association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr, taking into account the sky position, distance, and mass of the sources using a Bayesian criterion. Although the previous study has shown that GW190514 and GW190521 are a plausible hierarchical triple merger, their analysis is not significant with $p$-value of $\sim$0.14 [3] (generally, a $p$-value of $\leq$0.05 is considered ‘significant’). Here, we introduce a new approach and propose using multimessenger [4] signals to analyze the probability of the GW190514-GW190521 hierarchical triple merger. Our analysis (considering the primary masses) results greatly increase the probability, finding that a log Bayes factor of ${\rm log}\mathcal{B}=16.8~{}(21.5)$, corresponding to an odds ratio of $\mathcal{O}\sim$$1199~{}(2\times 10^{4}):$ 1, assuming an astrophysical prior odds of $\mathcal{P}\sim 10^{-5}$. This suggest strong evidence for the first hierarchical triple merger associated with an EM counterpart in the AGN formation channel. The finding of the multimessenger hierarchical triple merger in the AGN is crucial for understanding the formation mechanisms of massive BHs, the role of AGNs in hierarchical mergers, and the implications of multimessenger astronomy.
@@
We use the GW datasets released with GWTC-2.1 [2], labeled as “cosm” and “IMRPhenomXPHM.” This analysis was made possible following software packages: NumPy [62], SciPy [63], Matplotlib [64], emcee [65], IPython [66], corner [67], seaborn [68], and Astropy [69].